# Nearly forgotten results in development of physical cosmology

### v2

@@ 2 Start of GR studies in Russia
Georges Henri Joseph Édouard Lemaître (17 July 1894 – 20 June 1966) was a Belgian Catholic priest who made major contributions to cosmology. [^9] Lemaître was born in Charleroi, Belgium approximately 20 year before WWI. In 1911, Lemaître started to study engineering at the Catholic University of Louvain. In 1914, after the outbreak of World War I, Lemaître interrupted his studies and he entered the Belgian army as a volunteer. After the war, Lemaître abandoned engineering for the study of physics and mathematics. In the 1920s, Lemaître got scholarship and spent a few years in the USA, studied in Harvard and worked at the Harvard Observatory. In 1925 he came back to Belgium became a part-time lecturer at the Catholic University of Louvain. After a defence of his PhD thesis in MIT in 1927 he got a ordinary professor position in Louvain. Lemaître personally knew Vesto Slipher (for instance, it is known that he visited Slipher in Arizona and Hubble at Mount Wilson in the summer of 1925 [39]). Slipher analyzed Doppler spectral shifts and found that distant nebulae are preferably moving from us [40]. E. Hubble used Slipher’s data on velocities of distant nebulae, and now some historians of science proposed to rename the Hubble constant [41]. Interesting details on Lemaître’s life and scientific investigations are given in [42]. G. Lemaître played an exclusive role in a development of cosmological studies. As a recognition of his achievements at the Symposium devoted to 50 years since the introduction of the Primeval Atom concept an outstanding cosmologist James Peebles (he got Nobel prize in physics in 2019) wrote [43]: ”Physical scientists have a healthy attitude toward the history of their subject: by and large we ignore it. But it is good to pause now and then and consider the careers of those who through a combination of the right talent at the propitious time have had an exceptional influence on the progress of science. As I have noted on several occasions it seems to me that Georges Lemaître played a unique and remarkable role in setting out the program of research we now call physical cosmology”. Similar Friedmann’s results were obtained independently by Belgian Abbe George Lemaître, where a derivation of the law which describes velocities of distant objects as dependence on distances toward them was observationally discovered by E. Hubble, and later this law began to bear his name. While Friedmann was a skillful mathematician and he did not speculate about astronomical applications of his results, Lemaître had a deep knowledge in astronomy and always think about an opportunity to test theoretical laws with astronomical observations. So, in 1927 Lemaître published paper [44] in obscure Belgium journal where the expanding Universe solution of the Einstein equations and he analytically derived a relation between distance and velocity of distant galaxies (now it is called Hubble – Lemaître law). He obtained about 575 km/sec/Mpc [45, 46], not very different from Hubble’s ultimate value. From 24 to 29 October 1927, the 5th Solvay Congress on physics in Brussels was organized and Einstein participated at this activity. The main topic was ”Electrons and Photons” or in other words, discussion of main quantum mechanics laws which were discovered in 1925. H. Lorentz was the chair of the Congress (he was also the chair at the four previous Solvay Congresses on physics). Lemaître was not invited to those discussions since he was not famous as other participants of the Congress, [^10] Participants of the Vth Solvay Congress in physics are presented in https://en.wikipedia.org/wiki/Solvay_Conference#/media/File:Solvay_conference_1927.jpg. As the History Insider noted it is ”The Most Intelligent Photo Ever Taken” since 17 Nobel Prize Winners are presented at this photo. however, Louvain is only 20 km from the Belgium capital. Lemaître went to the Congress venue and during the Congress breaks Lemaître had conversations with Einstein about his cosmological solutions published in 1927. Einstein commented favourably the Lemaître’s mathematical skills but he rejected an expanding Universe concept. ”Your calculations are correct, but your physical insight is abominable,” concluded Einstein [47]. Einstein also directed Lemaître’s attention at the Friedmann’s cosmological studies. In spite of Einstein’s skepticism in respect to models of expanding Universe, Lemaître continued to work on interpretation since he knew Slipher’s results that majority of distant galaxies have redshifts [47, 48], it means that most galaxies are moving away from us. Using Cepheid variable stars (who are giant stars where there is dependence between their luminosities and periods) E. Hubble found a linear dependence between velocities of extragalactic sources and distances toward them [49], or in other words, $V=HR$, where $V$ is a velocity of extragalactic source, $R$ is a distance toward a chosen galaxy, $H$ is a constant, which is called now the Hubble constant (in this paper Hubble estimated the constant as 500 (km/s)/Mpc. [^11] It should be noted that the law of the linear dependence of velocity (or redshift) on distance was discovered by Hubble based on the motion of nearby galaxies that are not uniformly and isotropically located, as Lemaître assumed after Einstein, and before Hubble, he obtained an estimate of the proportionality parameter in this linear relationship, which is close in value to Hubble’s estimates. Many years after the discovery of the Hubble-Lemaître law (since it was proposed and accepted to call this dependence in such a way at the IAU congress in 2018 in Vienna), J. Peebles wrote ”This nearly homogeneous expansion is in striking contrast to the extremely clumpy space distribution” [50]. Currently, astronomers estimate this constant and get a value about an order of magnitude smaller, which is very unusual for physical constants, and the main problem and inaccuracies are related to measuring distances.
@@
As it was noted earlier, in 1930 Eddington proved that static cosmological solution proposed by Einstein is unstable and at the Meeting of Royal Astronomical Society on January 10, de Sitter expressed doubts in respect to Einstein’s model of static Universe taking into account Hubble’s results and at Meeting it was declared that it is necessary to search for non-stationary cosmological solutions and it was published about these ideas in February Issue of Observatory Journal. Usually Lemaître read the journal (in particular, this February issue) and sent a letter to Eddington [42] where he informed that he solved the problem 3 years ago: ”Dear Professor Eddington, I just read the February n° of the Observatory and your suggestion of investigating a nonstatistical intermediary solution between those of Einstein and de Sitter. I made these investigations two years ago. I consider a universe of curvature constant in space but increasing with time. And I emphasize the existence of a solution in which the motion of the nebulae is always a receding one…” Lemaître also attached copies of his paper [44] to the letter to Eddington. After reading the Lemaître’s letter Eddington understood that the problem to find a suitable cosmological solution which is consistent with astronomical observations has been solved and he started immediately to promote this breakthrough [22]. In particular, in a review [51] on a Silberstein’s book Eddington wrote ”Three years ago a very substantial advance in this subject was made by Abbé G. Lemaître…it renders obsolete the contest between Einstein’s and de Sitter’s cosmogonies… We can now prove that Einstein’s universe is unstable. The equilibrium having been disturbed, the universe will progress through a continuous series of intermediate states towards the limit represented by de Sitter’s universe. By Lemaître’s analysis the history of this progress can be studied; and the intermediate stages (one of which must represent the present state of the world) can be treated in detail.” Eddington was sure that is was found the solution that described a behavior of the Universe and he wrote [52]: ”We have recently learnt mainly through the work of the that this spherical space is expanding rather rapidly.” Eddington recommended to publish an English translation of Lemaître’s paper [44] in Monthly Notices of the Royal Astronomical Society (MNRAS) and after that William Smart, the Editor of MNRAS and RAS Secretary sent a letter to Lemaître requesting him to submit an English translation of his 1927 paper and to propose to him to become a member of this Society [42]. In Smart’s letter it was written that the translation should have not more than 72 paragraphs, and following Smart’s recommendations Lemaître reduced an article length and deleted the piece of text where the quantity (which called now the Hubble constant) was evaluated from available observational data on redshifts of moving galaxies. Some authors declared that it was a free Lemaître’s choice do not indicate in 1931 that in 1927 Lemaître discovered the law now named after Hubble [53] (since 2018 it is called the Hubble – Lemaître law) but other authors claimed that this decision was made under pressure from Smart [22] and for English speaking people Hubble had the absolute priority in discovering this law. In 1931 Lemaître translated his paper written in French [44] and published the English translation in MNRAS [54]. Also in 1931 Lemaître considered the birth of the Universe as a quantum physics phenomenon [55]. It was the first introduction of a hot cosmological model (which were called Big Bang following Hoyle’s ironical cliche). Later, these Lemaître’s ideas were transformed in the Primeval Atom hypothesis. It would interesting to note that Lemaître predicted a background radiation with a temperature around a few kelvins in his cosmological model of hot Universe. more precisely, in 1934 Lemaître wrote [56] ”…if all the atoms of the stars were equally distributed through space there would be about one atom per cubic yard, or the total energy would be that of an equilibrium radiation at the temperature of liquid hydrogen..” (see also comments in [57]) Lemaître hoped that his cosmological approach may explain an origin of cosmic rays (at this time Robert Millikan was interested in the subject and he proposed cosmic rays as a term for high enrgy particles from space). At the end of 1932, Millikan invited Lemaître to visit Caltech. On January 11, 1933, Lemaître was invited to deliver a lecture at the Mount Wilson Observatory, where E. Hubble worked. A. Einstein followed this Lemaître’s lecture. When journalists asked Einstein about his impression of Lemaître’s cosmological model, Einstein replied, “This is the most beautiful and satisfactory explanation of creation to which I have ever listened!” [^12] D. Lambert, Einstein and Lemaître : Two friends, two cosmologies…, https://inters.org/einstein-lemaitre. After that journalist Dunkan Aikman published interview with Lemaître with the title ”Lemaître follows two paths to truth: The famous scientist, who is also a priest, tells why he finds no conflict between science and religion”. This article was extremely popular over the world. Because the theses that religion can open the way to truth, and that religion does not contradict science, were unacceptable to Soviet ideologists. They found a way out of this situation, just as it was later done for genetics and cybernetics. Soviet ideologists came to the conclusion that the approach to constructing cosmological models based on general relativity should be declared pseudoscientific, and the Soviet scientists who worked on this topic should be declared pseudoscientists and admirers of Western bourgeois science. Thus, a division occurred in cosmology: Western cosmology, which considered realistic dynamic models of the Universe, and Soviet cosmology, which asserted that the Universe is infinite in time and space. It was done in contradiction to a famous Chekhov’s statement that there is no national science as there is no national table and Soviet ideologists declared that the Universe must be infinite in space and in time in contrast Friedmann – Lemaître point of view on the subject. Even if famous Soviet scientists considered non-stationary cosmological models describing expanding Universe in this case, they were rather strongly criticized by representatives of the Communist Party apparatus, and, unfortunately, some of the critics were quite good professionals in mathematics and physics and in spite of that they participated in actions on pressure on genuine science and its representatives. Let us consider one interesting case when Soviet scientific community decided to celebrate half-century since the creation of special relativity and the Division of Physics and Mathematics decided to organize a Session devoted to this event on November 30 – December 1, 1955. After the Session, on January 16, 1956 instructor of the Central Committee of the Communist Party of the Soviet Union (CC CPSU) A. S. Monin and the head of the Science Department of CC CPSU V. A. Kirillin sent a letter to the Presidium of the Soviet Academy of Sciences where they strongly criticized the Session where achievements of relativistic cosmology were discussed. The authors of the letter thought that the talks of E. M. Lifshits, Ya. B. Zeldovich, L. D. Landau and V. L. Ginzburg had significant disadvantages and were not criticized properly by the audience [58]. At the beginning of the letter, it was written, ”…The program of the session, prepared by the Organizing Committee, which included Academicians Tamm and Landau, Corresponding Member of the USSR Academy of Sciences Ginzburg and Professor Lifshitz, was unsatisfactory, as it included reports by Academician Landau, Corresponding Member of the USSR Academy of Sciences Ginzburg and Professor Lifshitz, who do not work in the field of relativity theory and are known for their nihilistic attitude to the development of methodological issues of this theory”. Time has passed, and now we know the importance of the criticized scientists for Soviet theoretical physics, and it is hard to believe that their colleagues criticized them genuinely, since both Monin and Kirillin later became great Soviet organizers of science and full academicians in the field of physics and its applications. Concerning E. M. Lifshits report at the session the letter authors wrote ”This report is devoted to a review of research on relativistic cosmology. He was a significant advocate of the idealistic ”theory of the expanding universe.” This ”theory”, which is an illegal extension of non-stationary solutions of Einstein’s gravitational equations to the universe as a whole, was built by Abbé Lemaître on the direct order of the Pope [the italic font in this phrase was chosen by AZ]. According to this ”theory,” the universe has a finite age; at the time of its formation, it occupied an infinitesimally small volume, and then began to expand; such an expansion is taking place at the present time.” In this letter the authors also criticized Tamm, who chaired the session where Lifshitz’s report was heard. Thus, the authors criticized and labeled as unprofessional almost all Soviet theoretical physicists who later became Nobel laureates, namely, I.E. Tamm, L.D. Landau, V.L. Ginzburg and only I.M. Frank was not criticised (probably because he was not presented at the Session). Naturally, the question arises as to the professionalism and impartiality of the experts rendering such a verdict on the incompetence of the most prominent Soviet theorists.
@@
Since 1948 G. Gamow started to work American project to create a superbomb or in other words, H-bomb. During this period, refined cross-section data on the nuclear interaction of light elements became known, and these data could be used to calculate helium nucleosynthesis in the initial minutes of the evolution of the universe. At the end of 1940s, G. Gamow proposed his version of the hot Universe model where he estimated a helium nucleosynthesis in primordial Universe [59, 60], and soon after that his students calculated Cosmic Microwave Background (CMB) radiation [61] which should have now temperature around 5 K. In consequent calculations the CMB temperature was slightly different, but always it was a few kelvins. It would be reasonable to note that similar estimates for background radion temperature were done by Lemaître as we noted earlier. The CMB (cosmic microwave background) radiation or relic radiation in Russian literature is one of the key signatures of the hot Universe model developed by G. Gamow in the 1940s and 1950s. The CMB radiation was discovered by T. Shmaonov at the Pulkovo Observatory several years before A. Penzias and R. Wilson (who were awarded the Nobel prize in 1978) [62]. [^13] For In 1940, Andrew McKellar, analyzing the spectra CN and CH, found that interstellar medium was very cold with a temperature of approximately 2 K, or more precisely 2.7, 2.1 and 0.8 K [63]. However, astronomers did not have a correct explanation for this phenomenon for decades. Now Shmaonov’s discovery is widely known not only in Russia but also in the world, for instance, there is a description of this Shmaonov’s result in a fundamental book [64] (10 years after publication of this book Peebles was awarded by Nobel prize in physics for CMB studies), and see also Trimble’s review [65]. T. A. Shmaonov presented his lecture on the discovery 60 years after the events in Fig. 3. In Pulkovo observatory Shmaonov’s supervisors were professor Naum Lvovich Kaidanovsky and professor Semyon Emmanuilovich Khaikin. Khaikin was one of the first L. I. Mandelstamm’s students in Moscow, a great expert in theoretical physics, non-linear physics, radiophysics, radio astronomy, and the dean of the Physics Department at Moscow State University. Khaikin was fired from Moscow State University, MEPhI, and Lebedev Physical Institute after being accused of promoting idealism and Machism. A nice book was recently published about this remarkable scientist and teacher [66]. Shmaonov asked Khaikin about possible interpretation of his discovery, and Khaikin replied that he had no explanation, but it has to be published because it may be very important [66]. Since 1930s until 1960s Soviet physicists did not cite Gamow’s papers at all, Khaikin worked as the head of the Radioastronomy Department at the Main Astronomical Observatory in Pulkovo, Gamow was one of the brightest representatives of Leningrad school of physics, and there is no doubt Gamow’s works were known to many leading Soviet physicists, such as Khaikin, but he preferred not to declare that his PhD student (Shmaonov) had received a confirmation of the Gamow’s predictions. Moreover, as indicated in the article on cosmology published in the Great Soviet Encyclopedia [67], scientists discussing models of the expanding universe are recognized as supporters of bourgeois and religious philosophy, which is hostile to the Soviet one. Earlier in the 1940s and 1950s, Khaikin had already been harshly criticized by his colleagues for allegedly promoting idealism and Machism at Moscow University and the Lebedev Physical Institute, which subsequently led to his dismissal. After defending his PhD thesis, T. Shmaonov left astronomy and later worked at the Institute of General Physics in the laser physics laboratory headed by A. M. Prokhorov, who got the Nobel prize in 1965 for the laser discovery and was Academician-Secretary of the Department of General Physics and Astronomy (1973–1993). In 1978, when A. Penzias and R. Wilson got their Nobel prize for CMB discovery, A. M. Prokhorov criticized Shmaonov that he did not inform properly scientific community about his CMB studies and did not promote his discovery. Sometimes people say that Shmaonov’s achievements were not known for many years, since they did not have an appropriate cosmological interpretation. There is a popular opinion that no one in the Soviet Union knew about the Gamow model of the hot Universe and the predictions of this model. However, it seams to me this interpretation is not correct. After all, generally speaking, not discussing something publicly does not mean that the subject of discussion is not known. The fact that the predictions of the hot universe model were not publicly discussed does not mean that these predictions were not known. So already in an article in 1963 in the jubilee collection dedicated to the 75th anniversary of Friedmann’s birth, Zeldovich critically discussed these predictions. As mentioned earlier, in the Soviet Union the models of expanding Universe (including Gamow’s ones) were considered inadequate descriptions of the Universe and their consideration was not welcomed by official ideology and philosophy. In addition, despite the fact that Gamow was one of the most famous Soviet theoretical physicists, he did not return from a scientific trip abroad without the permission of the authorities. Thus, the mention of Gamow’s works could be interpreted as a support for his disloyal attitude towards the Soviet government. A. D. Chernin called the story of the lack of a correct interpretation of Shmaonov’s results a sad episode and noted that an American historian of science had found and interpreted Shmaonov’s work in a cosmological context [68]. In another article dedicated to Gamow’s 90th birthday [69], Chernin showed how Gamow simply estimated the temperature of the cosmic microwave background radiation without a detailed analysis of primary nucleosynthesis.

### v3

@@ 2 Start of GR studies in Russia
Georges Henri Joseph Édouard Lemaître (17 July 1894 – 20 June 1966) was a Belgian Catholic priest who made major contributions to cosmology. [^8] Lemaître was born in Charleroi, Belgium approximately 20 year before WWI. In 1911, Lemaître started to study engineering at the Catholic University of Louvain. In 1914, after the outbreak of World War I, Lemaître interrupted his studies and he entered the Belgian army as a volunteer. After the war, Lemaître abandoned engineering for the study of physics and mathematics. In the 1920s, Lemaître got scholarship and spent a few years in the USA, studied in Harvard and worked at the Harvard Observatory. In 1925 he came back to Belgium became a part-time lecturer at the Catholic University of Louvain. After a defence of his PhD thesis in MIT in 1927 he got a ordinary professor position in Louvain. Lemaître personally knew Vesto Slipher (for instance, it is known that he visited Slipher in Arizona and Hubble at Mount Wilson in the summer of 1925 [42]). Slipher analyzed Doppler spectral shifts and found that distant nebulae are preferably moving from us [43]. E. Hubble used Slipher’s data on velocities of distant nebulae, and now some historians of science proposed to rename the Hubble constant [44]. Interesting details on Lemaître’s life and scientific investigations are given in [45]. G. Lemaître played an exclusive role in a development of cosmological studies. As a recognition of his achievements at the Symposium devoted to 50 years since the introduction of the Primeval Atom concept an outstanding cosmologist James Peebles (he got Nobel prize in physics in 2019) wrote [46]: ’’Physical scientists have a healthy attitude toward the history of their subject: by and large we ignore it. But it is good to pause now and then and consider the careers of those who through a combination of the right talent at the propitious time have had an exceptional influence on the progress of science. As I have noted on several occasions it seems to me that Georges Lemaître played a unique and remarkable role in setting out the program of research we now call physical cosmology’’. The results similar to Friedmann’s were independently obtained by the Belgian Abbe George Lemaître. In addition Lemaître derived of the law which describes velocities of distant objects as dependence on distances toward them was observationally discovered by E. Hubble, and later this law began to bear his name. While Friedmann was a skillful mathematician and he did not speculate about astronomical applications of his results, Lemaître had a deep knowledge in astronomy and always think about an opportunity to test theoretical laws with astronomical observations. So, in 1927 Lemaître published paper [47] in an obscure Belgium journal where the expanding Universe solution of the Einstein equations were found and he analytically derived a relation between distance and velocity of distant galaxies (now it is called Hubble – Lemaître law). He obtained about 575 km/sec/Mpc [48, 49], not very different from Hubble’s ultimate value. From 24 to 29 October 1927, the 5th Solvay Congress on physics in Brussels was organized and Einstein participated at this activity. The main topic was ’’Electrons and Photons’’ or in other words, discussion of main quantum mechanics laws which were discovered in 1925. H. Lorentz was the chair of the Congress (he was also the chair at the four previous Solvay Congresses on physics). Lemaître was not invited to those discussions since he was not so famous as other participants of the Congress, [^9] Participants of the Vth Solvay Congress in physics are presented in https://en.wikipedia.org/wiki/Solvay_Conference#/media/File:Solvay_conference_1927.jpg. As the History Insider noted it is ”The Most Intelligent Photo Ever Taken” since 17 Nobel Prize Winners are presented at this photo. however, Louvain is only 20 km from the Belgium capital. Lemaître went to the Congress venue and during the Congress breaks Lemaître had conversations with Einstein about his cosmological solutions published in 1927. Einstein commented favourably the Lemaître’s mathematical skills but he rejected an expanding Universe concept. ’’Your calculations are correct, but your physical insight is abominable,’’ concluded Einstein [50]. Einstein also directed Lemaître’s attention at the Friedmann’s cosmological studies. In spite of Einstein’s skepticism in respect to models of expanding Universe, Lemaître continued to work on interpretation since he knew Slipher’s results that majority of distant galaxies have redshifts [50, 51], it means that most galaxies are moving away from us. Using Cepheid variable stars (who are giant stars where there is dependence between their luminosities and periods) E. Hubble found a linear dependence between velocities of extragalactic sources and distances toward them [52], or in other words, $V=HR$, where $V$ is a velocity of extragalactic source, $R$ is a distance toward a chosen galaxy, $H$ is a constant, which is called now the Hubble constant (in this paper Hubble estimated the constant as 500 (km/s)/Mpc. [^10] It should be noted that the law of the linear dependence of velocity (or redshift) on distance was discovered by Hubble based on the motion of nearby galaxies that are not uniformly and isotropically located, as Lemaître assumed after Einstein, and before Hubble, he obtained an estimate of the proportionality parameter in this linear relationship, which is close in value to Hubble’s estimates. At the IAU congress in 2018 in Vienna it was proposed and accepted to call this dependence as the Hubble–Lemaître law. Many years after the discovery of the law J. Peebles wrote ”This nearly homogeneous expansion is in striking contrast to the extremely clumpy space distribution” [53]. Currently, astronomers estimate this constant and get a value about an order of magnitude smaller, which is very unusual for physical constants, and the main problem and inaccuracies are related to measuring distances.
@@
As it was noted earlier, in 1930 A. Eddington proved that the static cosmological solution proposed by Einstein is unstable and at the Meeting of Royal Astronomical Society on January 10, de Sitter expressed doubts in respect to the Einstein’s model of static Universe taking into account Hubble’s results and at Meeting it was declared that it is necessary to search for non-stationary cosmological solutions. These issues were published in February Issue of Observatory Journal. Usually Lemaître read the journal (in particular, this February issue) and sent a letter to Eddington [45] where he informed that he solved the problem 3 years ago: ’’Dear Professor Eddington, I just read the February n° of the Observatory and your suggestion of investigating a nonstatistical intermediary solution between those of Einstein and de Sitter. I made these investigations two years ago. I consider a universe of curvature constant in space but increasing with time. And I emphasize the existence of a solution in which the motion of the nebulae is always a receding one…’’ Lemaître also attached copies of his paper [47] to the letter to Eddington. After reading the Lemaître’s letter Eddington understood that the problem to find a suitable cosmological solution which is consistent with astronomical observations has been solved and he started immediately to promote this breakthrough [24]. In particular, in a review [54] on a Silberstein’s book Eddington wrote ’’Three years ago a very substantial advance in this subject was made by Abbé G. Lemaître…it renders obsolete the contest between Einstein’s and de Sitter’s cosmogonies… We can now prove that Einstein’s universe is unstable. The equilibrium having been disturbed, the universe will progress through a continuous series of intermediate states towards the limit represented by de Sitter’s universe. By Lemaître’s analysis the history of this progress can be studied; and the intermediate stages (one of which must represent the present state of the world) can be treated in detail.’’ Eddington was sure that is was found the solution that described a behavior of the Universe and he wrote [55]: ’’We have recently learnt mainly through the work of Prof. Lemaître that this spherical space is expanding rather rapidly.’’ Eddington recommended to publish an English translation of Lemaître’s paper [47] in Monthly Notices of the Royal Astronomical Society (MNRAS) and after that William Smart, the Editor of MNRAS and RAS Secretary sent a letter to Lemaître requesting him to submit an English translation of his 1927 paper and to propose to him to become a member of this Society [45]. In Smart’s letter it was written that the translation should have not more than 72 paragraphs, and following Smart’s recommendations Lemaître reduced an article length and deleted the piece of text where the quantity (which called now the Hubble constant) was evaluated from available observational data on redshifts of moving galaxies. Some authors declared that it was a free Lemaître’s choice to do not indicate in 1931 that in 1927 Lemaître discovered the law now named after Hubble [56] (since 2018 it is called the Hubble – Lemaître law). However, other authors claimed that this decision was made under pressure from Smart [24] and for English speaking people Hubble had the absolute priority in discovering this law. In 1931 Lemaître translated his paper written in French [47] and published the English translation in MNRAS [57]. Also in 1931 Lemaître considered the birth of the Universe as a quantum physics phenomenon [58]. It was the first introduction of a hot cosmological model (which were called Big Bang following Hoyle’s ironical cliche). Later, these Lemaître’s ideas were transformed in the Primeval Atom hypothesis. It would interesting to note that Lemaître predicted a background radiation with a temperature around a few kelvins in his cosmological model of hot Universe. more precisely, in 1934 Lemaître wrote [59] ’’…if all the atoms of the stars were equally distributed through space there would be about one atom per cubic yard, or the total energy would be that of an equilibrium radiation at the temperature of liquid hydrogen..’’ (see also comments in [60]). Lemaître hoped that his cosmological approach may explain an origin of cosmic rays (at this time Robert Millikan was interested in the subject and he proposed cosmic rays as a term for high enrgy particles from space). At the end of 1932, Millikan invited Lemaître to visit Caltech. On January 11, 1933, Lemaître was invited to deliver a lecture at the Mount Wilson Observatory, where E. Hubble worked. A. Einstein followed this Lemaître’s lecture. When journalists asked Einstein about his impression of Lemaître’s cosmological model, Einstein replied, “This is the most beautiful and satisfactory explanation of creation to which I have ever listened!”  [^11] D. Lambert, Einstein and Lemaître : Two friends, two cosmologies…, https://inters.org/einstein-lemaitre. . After that journalist Dunkan Aikman published interview with Lemaître with the title ’’Lemaître follows two paths to truth: The famous scientist, who is also a priest, tells why he finds no conflict between science and religion’’ (the interview was published in New York Times on February 19, 1933  [^12] https://inters.org/files/lemaitre_two_paths_truth_nyt1933.pdf.. This article was extremely popular over the world. Because the theses that religion can open the way to truth, and that religion does not contradict science, were unacceptable to Soviet ideologists. They found a way out of this situation, just as it was later done for genetics and cybernetics. Soviet ideologists came to the conclusion that the approach to constructing cosmological models based on general relativity should be declared pseudoscientific, and the Soviet scientists who worked on this topic should be declared pseudoscientists and admirers of Western bourgeois science. Thus, a division occurred in cosmology: Western cosmology, which considered realistic dynamic models of the Universe, and Soviet cosmology, which asserted that the Universe is infinite in time and space. It was done in contradiction to a famous Chekhov’s statement that there is no national science as there is no national multiplication table. So, Soviet ideologists declared that the Universe must be infinite in space and in time in contrast Friedmann – Lemaître point of view on the subject.
Even if famous Soviet scientists considered non-stationary cosmological models describing expanding Universe in this case, they were rather strongly criticized by representatives of the Communist Party apparatus, and, unfortunately, some of the critics were quite good professionals in mathematics and physics and in spite of that they participated in actions on pressure on genuine science and its representatives. Let us consider one interesting case when Soviet scientific community decided to celebrate half-century since the creation of special relativity and the Division of Physics and Mathematics decided to organize a Session devoted to this event on November 30 – December 1, 1955. After the Session, on January 16, 1956 instructor of the Central Committee of the Communist Party of the Soviet Union (CC CPSU) A. S. Monin [^13] https://en.wikipedia.org/wiki/Andrei_Monin. and the head of the Science Department of CC CPSU V. A. Kirillin [^14] https://en.wikipedia.org/wiki/Vladimir_Kirillin. sent a letter to the Presidium of the Soviet Academy of Sciences where they strongly criticized the Session where achievements of relativistic cosmology were discussed. The authors of the letter declared that the talks of E. M. Lifshits, Ya. B. Zeldovich, L. D. Landau and V. L. Ginzburg had significant disadvantages and were not criticized properly by the audience [61]. At the beginning of the letter, it was written, ’’…The program of the session, prepared by the Organizing Committee, which included Academicians Tamm and Landau, Corresponding Member of the USSR Academy of Sciences Ginzburg and Professor Lifshitz, was unsatisfactory, as it included reports by Academician Landau, Corresponding Member of the USSR Academy of Sciences Ginzburg and Professor Lifshitz, who do not work in the field of relativity theory and are known for their nihilistic attitude to the development of methodological issues of this theory’’. Time has passed, and now we know the importance of the criticized scientists for Soviet theoretical physics, and it is hard to believe that their colleagues criticized them genuinely, since both Monin and Kirillin later became great Soviet organizers of science and full academicians in the field of physics and its applications. Concerning E. M. Lifshits report at the session the letter authors wrote ’’This report is devoted to a review of research on relativistic cosmology. He was a significant advocate of the idealistic "theory of the expanding universe." This "theory", which is an illegal extension of non-stationary solutions of Einstein’s gravitational equations to the universe as a whole, was built by Abbé Lemaître on the direct order of the Pope [the italic font in this phrase was chosen by AZ]. According to this "theory," the universe has a finite age; at the time of its formation, it occupied an infinitesimally small volume, and then began to expand; such an expansion is taking place at the present time.’’ In this letter the authors also criticized Tamm, who chaired the session where Lifshitz’s report was heard. Thus, the authors criticized and labeled as unprofessional almost all Soviet theoretical physicists who later became Nobel laureates, namely, I.E. Tamm, L.D. Landau, V.L. Ginzburg and only I.M. Frank was not criticised (probably because he was not presented at the Session). Naturally, the question arises as to the professionalism and impartiality of the experts rendering such a verdict on the incompetence of the most prominent Soviet theorists.
@@
Since 1948 G. Gamow started to work American project to create a superbomb or in other words, H-bomb. During this period, refined cross-section data on the nuclear interaction of light elements became known, and these data could be used to calculate helium nucleosynthesis in the initial minutes of the evolution of the universe. At the end of 1940s, G. Gamow proposed his version of the hot Universe model where he estimated a helium nucleosynthesis in primordial Universe [62, 63], and soon after that his students calculated Cosmic Microwave Background (CMB) radiation [64] which should have now temperature around 5 K. In consequent calculations the CMB temperature was slightly different, but always it was a few kelvins. It would be reasonable to note that similar estimates for background radion temperature were done by Lemaître as we noted earlier. The CMB (cosmic microwave background) radiation or relic radiation in Russian literature is one of the key signatures of the hot Universe model developed by G. Gamow in the 1940s and 1950s. The CMB radiation was discovered by T. Shmaonov at the Pulkovo Observatory several years before A. Penzias and R. Wilson (who were awarded the Nobel prize in 1978) [65]. [^15] In 1940, Andrew McKellar, analyzing the spectra CN and CH, found that interstellar medium was very cold with a temperature of approximately 2 K, or more precisely 2.7, 2.1 and 0.8 K [66]. However, astronomers did not have a correct explanation for this phenomenon for decades. Now Shmaonov’s discovery is widely known not only in Russia but also in the world, for instance, there is a description of this Shmaonov’s result in a fundamental book [67] (10 years after publication of this book Peebles was awarded by Nobel prize in physics for CMB studies), and see also Trimble’s review [68]. T. A. Shmaonov presented his lecture on the discovery 60 years after the events, see Fig. 3. In Pulkovo observatory Shmaonov’s supervisors were professor Naum Lvovich Kaidanovsky and professor Semyon Emmanuilovich Khaikin. Khaikin was one of the first L. I. Mandelstamm’s students in Moscow, a great expert in theoretical physics, non-linear physics, radiophysics, radio astronomy, and the dean of the Physics Department at Moscow State University. Khaikin was fired from Moscow State University, MEPhI, and Lebedev Physical Institute after being accused of promoting idealism and Machism. A nice book was recently published about this remarkable scientist and teacher [69]. When Shmaonov obtained his observational results he asked Khaikin about possible interpretation of his discovery, and Khaikin replied that he had no explanation, but he recommended to published them because it may be very important [69]. Since 1930s until 1960s Soviet physicists did not cite Gamow’s papers at all, Khaikin worked as the head of the Radioastronomy Department at the Main Astronomical Observatory in Pulkovo, Gamow was one of the brightest representatives of Leningrad school of physics, and there is no doubt Gamow’s works were known to many leading Soviet physicists, such as Khaikin, but he preferred not to declare that his PhD student (Shmaonov) had received a confirmation of the Gamow’s predictions. Moreover, as indicated in the article on cosmology published in the Great Soviet Encyclopedia [70], scientists discussing models of the expanding universe are recognized as supporters of bourgeois and religious philosophy, which is hostile to the Soviet one. Earlier in the 1940s and 1950s, Khaikin had already been harshly criticized by his colleagues for allegedly promoting idealism and Machism at Moscow University and the Lebedev Physical Institute, which subsequently led to his dismissal. After defending his PhD thesis, T. Shmaonov left astronomy and later worked at the Institute of General Physics in the laser physics laboratory headed by A. M. Prokhorov, who got the Nobel prize in 1965 for the laser discovery and was Academician-Secretary of the Department of General Physics and Astronomy (1973–1993). In 1978, when A. Penzias and R. Wilson got their Nobel prize for CMB discovery, A. M. Prokhorov criticized Shmaonov that he did not inform properly scientific community about his CMB studies and did not promote his discovery. Sometimes people say that Shmaonov’s achievements were not known for many years, since they did not have an appropriate cosmological interpretation. There is a popular opinion that no one in the Soviet Union knew about the Gamow model of the hot Universe and the predictions of this model. However, it seams to me this interpretation is not correct. After all, generally speaking, not discussing something publicly does not mean that the subject of discussion is not known. The fact that the predictions of the hot universe model were not publicly discussed does not mean that these predictions were not known. So already in an article in 1963 in the jubilee collection dedicated to the 75th anniversary of Friedmann’s birth, Zeldovich critically discussed these predictions done by G. Gamow. As mentioned earlier, in the Soviet Union the models of expanding Universe (including Gamow’s ones) were considered inadequate descriptions of the Universe and their consideration was not welcomed by official ideology and philosophy. In addition, despite the fact that Gamow was one of the most famous Soviet theoretical physicists, he did not return from a scientific trip abroad without the permission of the authorities. Thus, the mention of Gamow’s works could be interpreted as a support for his disloyal attitude towards the Soviet government. A. D. Chernin called the story of the lack of a correct interpretation of Shmaonov’s results a sad episode and noted that an American historian of science had found and interpreted Shmaonov’s work in a cosmological context [71]. In another article dedicated to Gamow’s 90th birthday [72], Chernin showed how Gamow simply estimated the temperature of the cosmic microwave background radiation without a detailed analysis of primary nucleosynthesis.